# FastGS: Training 3D Gaussian Splatting in 100 Seconds

### v1

@@ 4.3 Multi-view Consistent Pruning
The vanilla 3DGS [kerbl3Dgaussians] removes Gaussians with low opacity or overly large scale but cannot effectively address redundancy. Recent pruning strategies [ali2024trimming, fan2024lightgaussian, fang2024mini] similarly fail to eliminate redundancy and can even significantly degrade rendering quality. In all cases, they do not determine Gaussian redundancy based on multi-view consistency. As illustrated in Fig. 3, Speedy-Splat [hanson2025speedy] computes the pruning score by accumulating Gaussian-associated Hessian approximations across all training views rather than capturing each Gaussian’s true contribution to multi-view rendering quality. Hence, it leads to degraded rendering quality due to its indirect use of multi-view consistency, as visualized on the right of Fig. 3. To remove truly redundant Gaussians, we propose a new, simple pruning strategy VCP based on multi-view consistency. As illustrated in Fig. 3, similar to VCD, it evaluates the score according to each Gaussian’s impact on multi-view reconstruction quality. As shown on the right of Fig. 3, VCP removes a significant number of redundant Gaussians while preserving rendering quality. We then detail how VCP is implemented.
@@
$s_{i}^{-}=\mathcal{N}\left(\sum_{j=1}^{K}{\left(\sum_{p}^{\varOmega_{i}}{\mathbb{I}\left(\mathcal{M}_{mask}^{j}\left(p\right)=1\right)}\right)\cdot E_{\mathrm{photo}}^{j}}\right),$ (12)
where $\mathcal{N}(\cdot)$ denotes a min–max normalization function. Here, $s_{i}^{-}$ can be interpreted as a quantitative measure of the contribution of Gaussian primitive $\mathcal{G}_{i}$ to the degradation of the overall rendering quality. A Gaussian primitive $\mathcal{G}_{i}$ is selected for pruning if its importance score $s^{-}_{i}$ exceeds a predefined threshold $\tau_{-}$, indicating that it has relatively low contribution to rendering quality across multiple views.
@@ 4.4 Compact Box
During the preprocessing stage of rasterization, the vanilla 3DGS [kerbl3Dgaussians] uses the 3-sigma rule to obtain 2D ellipses, generating many Gaussian–tile pairs that introduce computational redundancy and reduce rendering efficiency. Speedy-Splat [hanson2025speedy] partially addresses this with precise tile intersect, yet we observe that some 2D Gaussians still have negligible impact on pixels in certain tiles. To further reduce unnecessary pairs, we propose Compact Box, which prunes Gaussian–tile pairs with minimal contribution based on their Mahalanobis distance from the Gaussian center, thereby accelerating rendering while maintaining quality. Details are provided in the supplementary material.
@@ 4.5 Optimization
Table 1: Quantitative comparisons with existing 3DGS fast optimization methods. With FastGS, the training of 3DGS can be completed within 100 seconds on average, while achieving comparable rendering quality to the other methods. Best results are marked as best score, second best score and third best score. Time is reported in minutes.
[…]
3DGS [kerbl3Dgaussians] 20.93 27.53 0.813 0.221 2.63M 146 19.77 29.71 0.903 0.241 2.46M 158 11.34 23.71 0.850 0.170 1.57M 195
3DGS-accel [kerbl3Dgaussians, mallick2024taming] 10.94 27.52 0.813 0.221 2.63M 188 8.87 29.74 0.903 0.242 2.46M 230 6.96 23.85 0.850 0.170 1.57M 252
Mini-Splatting [fang2024mini] 17.71 27.33 0.821 0.217 0.53M 564 13.35 29.99 0.907 0.244 0.56M 624 9.06 23.46 0.844 0.181 0.30M 756
DashGaussian [chen2025dashgaussian] 6.38 27.73 0.817 0.217 2.40M 154 4.16 29.65 0.906 0.246 1.94M 208 4.28 24.00 0.853 0.178 1.21M 240
FastGS-20k (Ours) 1.32 27.35 0.793 0.268 0.39M 568 0.88 29.93 0.901 0.274 0.22M 687 0.95 23.81 0.835 0.216 0.25M 632
FastGS (Ours) 1.92 27.54 0.796 0.264 0.38M 584 1.28 30.03 0.901 0.270 0.22M 714 1.32 24.15 0.839 0.210 0.24M 655
FastGS-Big (Ours) 3.59 27.93 0.820 0.216 1.16M 472 2.00 30.12 0.907 0.242 0.65M 607 2.03 24.39 0.855 0.175 0.54M 569
Figure 4: Qualitative results of Tab. 1. We present qualitative results on the “kitchen” and “stump” scenes from the Mip-NeRF 360 dataset [barron2022mip], the “playroom” scene from the Deep Blending dataset [deng2022depth], and the “truck” scene from the Tanks&Temples dataset [knapitsch2017tanks].
@@ 5.1 Experimental Setup
Datasets. Same as vanilla 3DGS [kerbl3Dgaussians], we conduct experiments on three real-world datasets: Mip-NeRF 360 [barron2022mip], Deep-Blending [hedman2018deep], and Tanks & Temples [knapitsch2017tanks]. For other challenging tasks, we evaluate dynamic scene reconstruction on D-NeRF [pumarola2021d], NeRF-DS [yan2023nerf], and Neu3D [li2022neural] datasets. Tanks & Temples [knapitsch2017tanks], LLFF [mildenhall2019local], BungeeNeRF [xiangli2022bungeenerf] and Replica [replica19arxiv] are respectively used for surface reconstruction, sparse-view reconstruction, large-scale reconstruction and SLAM.
[…]
3DGS [kerbl3Dgaussians] 20.93 27.53 0.813 0.221 2.63M 146 19.77 29.71 0.903 0.241 2.46M 158 11.34 23.71 0.850 0.170 1.57M 195
+Ours 9.18 27.54 0.796 0.264 0.39M 504 9.60 30.09 0.901 0.270 0.22M 618 4.93 24.24 0.840 0.209 0.24M 586
Mip-Splatting [yu2024mip] 26.20 27.89 0.837 0.176 3.98M 224 23.94 29.35 0.899 0.241 3.48M 219 14.57 23.77 0.856 0.158 2.36M 300
+Ours 2.98 27.95 0.828 0.208 0.83M 606 1.74 29.68 0.899 0.274 0.20M 698 2.01 24.18 0.843 0.200 0.36M 729
Scaffold-GS [lu2024scaffold] 18.37 27.70 0.812 0.226 0.57M 194 13.31 30.09 0.905 0.256 0.18M 307 9.83 24.09 0.851 0.175 0.26M 261
+Ours 5.06 27.68 0.809 0.220 0.30M 281 2.82 30.00 0.900 0.267 0.08M 423 3.77 24.15 0.849 0.180 0.14M 332
3DGS-accel [kerbl3Dgaussians, mallick2024taming] 10.94 27.52 0.813 0.221 2.63M 188 8.87 29.74 0.903 0.242 2.46M 230 6.96 23.85 0.850 0.170 1.57M 252
+Ours 1.92 27.54 0.796 0.264 0.38M 584 1.28 30.03 0.901 0.270 0.22M 714 1.32 24.15 0.839 0.210 0.24M 655
@@
NeRF-DS [yan2023nerf] Deformable-3DGS [yang2024deformable] 7.86 23.83 0.851 0.180 0.13M 103
+Ours 2.75 23.90 0.854 0.176 0.03M 484
Neu3D [li2022neural] Deformable-3DGS [yang2024deformable] 25.11 27.78 0.902 0.158 0.20M 73
+Ours 8.70 30.25 0.919 0.140 0.08M 172
Implementation Details. All methods, including ours and the other compared approaches, are trained for 30K iterations using the Adam [kingma2014adam] optimizer. For our approach, we set $K=10$ and $\lambda=0.2$ in all experiments. In the base setting, densification is performed every 500 iterations until the 15,000-th iteration. To ensure fairness, all experiments are conducted with an NVIDIA RTX 4090 GPU, and all comparison methods are implemented using their official code. The default configuration of FastGS is 3DGS-accel [kerbl3Dgaussians, mallick2024taming] with the proposed VCD, VCP, and CB. To achieve extreme training acceleration, the rendering quality of our method is not the highest. Therefore, we provide a variant, FastGS_\_Big, which achieves both the highest rendering quality and the fastest training speed, where densification is executed once every 100 iterations. Further details are in the supplementary material.
Figure 5: Visualization results on four representative tasks. We present the rendered results of the “coffee-martini”, “Ignatius”, “fern”, and “bilbao” scenes from the Neu3D [li2022neural], Tanks and Temples [knapitsch2017tanks], LLFF [mildenhall2019local], and BungeeNeRF [xiangli2022bungeenerf] datasets, respectively.
@@ 5.2 Comparison with Fast Optimization Methods
Baselines. We adopt 3DGS-accel [kerbl3Dgaussians, mallick2024taming] as our baseline which preserves the original 3DGS [kerbl3Dgaussians] pipeline and incorporates the efficient per-splat parallel backpropagation proposed by Taming-3DGS [mallick2024taming]. Beyond this baseline, we further conduct comparisons against several recent SOTA acceleration techniques, including DashGaussian [chen2025dashgaussian], Mini-Splatting [fang2024mini], Speedy-Splat [hanson2025speedy], and Taming-3DGS [mallick2024taming], which represent complementary advances in accelerating training from different perspectives.
Quantitative Results. As shown in Tab. 1, FastGS achieves the fastest training with comparable rendering quality. On average, a scene can be trained within 100 seconds, and the fastest case takes only 77 seconds. This benefit comes from our effective multi-view consistent densification and pruning strategies. Speedy-Splat [hanson2025speedy] fails to fully consider multi-view consistency during pruning, leading to a significant drop in rendering quality. Taming3DGS [mallick2024taming] applies weak multi-view consistency constraints, resulting in excessive Gaussian growth and slower training. The current SOTA, DashGaussian [chen2025dashgaussian], achieves high rendering quality yet lacks an effective pruning strategy to remove redundant Gaussians. Its scene optimization still retains several million Gaussians, limiting training speed. In contrast, our variant FastGS-Big surpasses DashGaussian [chen2025dashgaussian] by more than 0.2 dB in rendering quality, reduces training time by over 43.7%, and cuts the number of Gaussians by more than half. These results demonstrate the superiority of our multi-view consistent densification and pruning strategies.
We visualize the results of Tab. 1 in Fig. 4, providing a detailed comparison of reconstructed scenes across different methods. Our method FastGS achieves comparable rendering quality without losing any details. In addition, our variant FastGS-Big reconstructs scenes more accurately, with more precise geometric structures under the multi-view consistency constraint.
@@
Tanks & Temples [knapitsch2017tanks] PGSR [chen2024pgsr] 32.92 24.20 0.857 0.149 1.56M 87 0.57
+Ours 15.73 24.37 0.845 0.190 0.40M 210 0.55
Mip-NeRF 360 [barron2022mip] PGSR [chen2024pgsr] 74.70 27.22 0.832 0.183 3.79M 45 -
+Ours 11.68 27.23 0.813 0.240 0.49M 202 -
5.3 Generality of FastGS
Baselines. Deformable-3DGS [yang2024deformable], PGSR [chen2024pgsr], DropGaussian [park2025dropgaussian], OctreeGS [ren2024octree] and Photo-SLAM [hhuang2024photoslam] are selected as the backbones for dynamic scene reconstruction, surface reconstruction, sparse-view reconstruction, large-scale reconstruction and SLAM, respectively.
@@ 5.3 Generality of FastGS
Enhancing Various Tasks. We further test several SOTA methods combined with our framework across these tasks. As shown in Tab. 3 and Tab. 4, our method improves the training speed of all baselines by 2-7× while preserving rendering quality. We visualize some results in Fig. 5. This improvement demonstrates the strong generality of our approach. We argue that this benefits from the multi-view consistency, which is fundamental to various reconstruction tasks. More results are provided in the supplementary material and we will open-source our code.
@@ 5.4 Ablation Study
We adopt 3DGS-accel [kerbl3Dgaussians, mallick2024taming] as our baseline and systematically evaluate the contribution of each proposed module. By adding each component individually, we analyze its impact on both reconstruction quality and training efficiency, providing insights into the effectiveness of our design.
Multi-View Consistent Densification. We first evaluate the effect of the densification strategy VCD. As shown in  Tab. 5, VCD achieves nearly 3× faster training without any loss in reconstruction quality. This is because our Gaussian addition is guided by stricter multi-view consistency, which prevents the addition of redundant Gaussians.  Tab. 5 further validates this by showing that with VCD, the number of Gaussians is reduced by 80%. This ablation study demonstrates the effectiveness of VCD for accelerating training.
@@
3DGS-accel 10.94 27.52 0.813 0.221 2.63M 188
+VCD. 3.60 27.58 0.798 0.260 0.51M 228
+VCP. 5.60 27.53 0.812 0.228 1.87M 292
+CB. 10.08 27.51 0.810 0.219 2.76M 325
Full 1.92 27.54 0.796 0.264 0.38M 584
Multi-View Consistent Pruning. Next, we evaluate the effectiveness of VCP. As shown in  Tab. 5, adding VCP shortens the training time by 48.8% and reduces the number of Gaussians by 28.9%, without sacrificing rendering quality. This is because our method effectively removes redundant Gaussians while preserving those critical for scene reconstruction. The strict multi-view consistency evaluation for each deleted Gaussian ensures this effectiveness, demonstrating that VCP is highly effective.
Compact Box. Finally, we evaluate the effectiveness of Compact Box. As shown in  Tab. 5, adding CB shortens the training time by 7.8% while achieving the comparable rendering quality. This demonstrates that CB can accelerate training without degrading reconstruction quality.

### v2

@@ 4.3 Multi-view Consistent Pruning
The vanilla 3DGS [kerbl3Dgaussians] removes Gaussians with low opacity or overly large scale, but cannot effectively address redundancy. Recent pruning strategies [ali2024trimming, fan2024lightgaussian, fang2024mini, salman2024elmgs, girish2024eagles] similarly fail to eliminate redundancy and can even significantly degrade rendering quality. In all cases, they do not determine Gaussian redundancy based on multi-view consistency. As illustrated in Fig. 3, Speedy-Splat [hanson2025speedy] computes the pruning score by accumulating Gaussian-associated Hessian approximations across all training views. Hence, it leads to degraded rendering quality due to its indirect use of multi-view consistency, as visualized on the right of Fig. 3. To remove truly redundant Gaussians, we propose a new, simple pruning strategy VCP based on multi-view consistency, as illustrated in Fig. 3. Similar to VCD, it evaluates the score according to each Gaussian’s impact on multi-view reconstruction quality. As shown on the right of Fig. 3, VCP removes a significant number of redundant Gaussians while preserving rendering quality. We then detail how VCP is implemented.
@@
$s_{p}^{i}=\mathcal{N}\left(\sum_{j=1}^{K}{\left(\sum_{p}^{\varOmega_{i}}{\mathbb{I}\left(\mathcal{M}_{mask}^{j}\left(p\right)=1\right)}\right)\cdot E_{\mathrm{photo}}^{j}}\right).$ (11)
Here, $s_{p}^{i}$ can be interpreted as a quantitative measure of the contribution of the Gaussian primitive $\mathcal{G}_{i}$ to the degradation of the overall rendering quality. A Gaussian primitive $\mathcal{G}_{i}$ is selected for pruning if its pruning score $s^{i}_{p}$ exceeds a predefined threshold $\tau_{p}$, indicating that it has relatively low contribution to rendering quality across multiple views.
Figure 4: Compact box. Compared with vanilla 3DGS [kerbl3Dgaussians] and Speedy-Splat [hanson2025speedy], incorporating CB leads to a reduced number of Gaussian-tile pairs.
[…]
Figure 5: Qualitative results of Tab. 1. We present qualitative results on the kitchen and stump scenes from Mip-NeRF 360 [barron2022mip], the playroom scene from Deep Blending [deng2022depth], and the truck scene from Tanks & Temples [knapitsch2017tanks]. Notably, the rendered results of the kitchen scene under multiple viewpoints demonstrate that our method achieves more consistent details across views.
@@ 4.4 Compact Box
During the preprocessing stage of rasterization, the vanilla 3DGS [kerbl3Dgaussians] uses the 3-sigma rule to obtain 2D ellipses, generating many Gaussian–tile pairs that introduce computational redundancy and reduce rendering efficiency. Speedy-Splat [hanson2025speedy] partially addresses this with precise tile intersection, yet we observe that some 2D Gaussians still have a negligible impact on pixels in certain tiles. As illustrated in Fig. 4, to further reduce unnecessary pairs, we propose compact box, which prunes Gaussian–tile pairs with minimal contribution based on Mahalanobis distance from the Gaussian center, thereby accelerating rendering while maintaining quality. Details are provided in Sec. 8 of the supplementary material.
Table 1: Quantitative comparisons with existing 3DGS fast optimization methods. With FastGS, the training of 3DGS can be completed in around 100 seconds, while achieving comparable rendering quality to the other methods. Best results are marked as best score, second best score, and third best score. Time is reported in minutes.
[…]
3DGS [kerbl3Dgaussians] 20.93 27.53 0.812 0.221 2.63M 146 19.77 29.71 0.903 0.241 2.46M 158 11.34 23.71 0.850 0.170 1.57M 195
Mini-Splatting [fang2024mini] 17.69 27.32 0.821 0.217 0.53M 567 13.35 29.99 0.907 0.244 0.56M 624 9.06 23.46 0.844 0.181 0.30M 756
DashGaussian [chen2025dashgaussian] 6.35 27.73 0.817 0.218 2.40M 155 4.16 29.65 0.906 0.246 1.94M 208 4.28 24.00 0.853 0.178 1.21M 240
FastGS (Ours) 1.93 27.56 0.797 0.261 0.40M 579 1.28 30.03 0.901 0.270 0.22M 714 1.32 24.15 0.839 0.210 0.24M 655
FastGS-Big (Ours) 3.58 27.93 0.820 0.216 1.15M 469 2.00 30.12 0.907 0.243 0.65M 607 2.03 24.39 0.855 0.175 0.54M 569
@@ 5.1 Experimental Setup
Datasets. Same as vanilla 3DGS [kerbl3Dgaussians], we conduct experiments on three real-world datasets: Mip-NeRF 360 [barron2022mip], Deep-Blending [hedman2018deep], and Tanks & Temples [knapitsch2017tanks]. Moreover, we evaluate dynamic scene reconstruction on D-NeRF [pumarola2021d], NeRF-DS [yan2023nerf], and Neu3D [li2022neural] datasets. Tanks & Temples [knapitsch2017tanks], LLFF [mildenhall2019local], BungeeNeRF [xiangli2022bungeenerf], and Replica [replica19arxiv] are respectively used for surface reconstruction, sparse-view reconstruction, large-scale reconstruction, and SLAM.
Figure 6: Visualization results on four representative tasks. We present the rendered results of the coffee-martini, fern, Ignatius, and bilbao scenes from the Neu3D [li2022neural], LLFF [mildenhall2019local], Tanks & Temples [knapitsch2017tanks], and BungeeNeRF [xiangli2022bungeenerf] datasets, respectively.
@@
NeRF- Deformable-3DGS [yang2024deformable] 7.86 23.83 0.851 0.180 0.13M 103
DS [yan2023nerf] +Ours 2.75 23.90 0.854 0.176 0.03M 484
Neu3D Deformable-3DGS [yang2024deformable] 26.16 26.74 0.888 0.168 0.21M 69
[li2022neural] +Ours 9.29 29.29 0.908 0.146 0.09M 161
Implementation Details. All methods, including ours and the other compared approaches, are trained for 30K iterations using the Adam [kingma2014adam] optimizer. For our approach, we set $K=10$ and $\lambda=0.2$ in all experiments. In the base setting, densification is performed every 500 iterations until the 15,000th iteration, and pruning is executed every 500 iterations before 15K and every 3,000 iterations afterwards. To ensure fairness, all experiments are conducted with an NVIDIA RTX 4090 GPU, and all comparison methods are implemented using their official code. The default configuration of FastGS builds upon 3DGS-accel [kerbl3Dgaussians, mallick2024taming], detailed in Tab. 13, and incorporates the proposed VCD, VCP, and CB. To achieve extreme training acceleration, the rendering quality of our method is not the highest. Therefore, we provide a variant, FastGS-Big, which achieves both the highest rendering quality and the fastest training speed, where densification is executed once every 100 iterations. Further details are in Sec. 9.1 of the supplementary material.
@@ 5.2 Comparison with Fast Optimization Methods
Baselines. We report comparisons with the SOTA fast optimization methods, including DashGaussian [chen2025dashgaussian], Mini-Splatting [fang2024mini], Speedy-Splat [hanson2025speedy], and Taming-3DGS [mallick2024taming], together with the vanilla 3DGS [kerbl3Dgaussians] for reference. These methods represent complementary approaches to accelerating training from different perspectives.
Table 3: Quantitative results of sparse-view reconstruction. Our method achieves an average 2.56× training speed-up.
Method LLFF [mildenhall2019local] (9-view)
[…]
DropGaussian [park2025dropgaussian] 1.41 26.13 0.874 0.089 0.42M 154
+Ours 0.55 26.14 0.873 0.095 0.21M 189
Quantitative Results. As shown in Tab. 1 and Fig. 5, FastGS achieves the fastest training with comparable rendering quality. A scene can be trained in around 100 seconds, and the fastest case takes only 77 seconds. Taming-3DGS [mallick2024taming] applies weak multi-view consistency constraints, resulting in excessive Gaussians and slower training. Similarly, the pruning strategy of Speedy-Splat [hanson2025speedy] leads to a significant drop in rendering quality. The current SOTA, DashGaussian [chen2025dashgaussian], achieves high rendering quality. However, its scene optimization still retains several million Gaussians, which limits the training speed. In contrast, our variant FastGS-Big surpasses DashGaussian [chen2025dashgaussian] by more than 0.2 dB in rendering quality, reduces training time by 43.6%, and cuts the number of Gaussians by more than half. These results demonstrate the superiority of our multi-view consistent densification and pruning strategies.
5.3 Generality of FastGS
Baselines. Deformable-3DGS [yang2024deformable], PGSR [chen2024pgsr], DropGaussian [park2025dropgaussian], OctreeGS [ren2024octree], and Photo-SLAM [hhuang2024photoslam] are selected as the backbones for dynamic scene reconstruction, surface reconstruction, sparse-view reconstruction, large-scale reconstruction, and SLAM, respectively.
Enhancing Various Tasks. We further test several SOTA methods combined with our framework across these tasks. As shown in Tab. 2, Tab. 3, Tab. 4, and Tab. 5, our method improves the training speed of all baselines by 2-6× while preserving rendering quality. We visualize rendered results in Fig. 6, there is no degradation in rendering quality across multiple tasks. This improvement demonstrates the strong generality of our approach. We argue that this benefits from the multi-view consistency, which is fundamental to various reconstruction tasks. More details and results are provided in Sec. 9.2 of the supplementary material.
@@ 5.3 Generality of FastGS
Tanks & PGSR [chen2024pgsr] 32.28 24.20 0.857 0.149 1.56M 87 0.57
Temples [knapitsch2017tanks] +Ours 15.74 24.37 0.845 0.190 0.40M 210 0.55
Mip-NeRF PGSR [chen2024pgsr] 74.70 27.22 0.832 0.183 3.79M 45 -
360 [barron2022mip] +Ours 11.68 27.23 0.813 0.240 0.49M 202 -
[…]
Method BungeeNeRF [xiangli2022bungeenerf]
[…]
Octree-GS [ren2024octree] 21.18 28.04 0.917 0.093 0.99M 141
+Ours 9.68 28.04 0.910 0.102 0.74M 162
@@ 5.4 Ablation Study
We adopt 3DGS-accel [kerbl3Dgaussians, mallick2024taming] as our baseline, which preserves the vanilla 3DGS [kerbl3Dgaussians] pipeline and integrates per-splat parallel backpropagation and accelerated SH optimization from Taming-3DGS [mallick2024taming], along with optimizer scheduling. We then systematically evaluate the contribution of each proposed module based on this baseline. By adding each component individually, we analyze its impact on both reconstruction quality and training efficiency.
Multi-View Consistent Densification. We first evaluate the effect of the densification strategy VCD. As shown in  Tab. 7, VCD achieves over 2× faster training without any loss in reconstruction quality. This is because our Gaussian addition is guided by stricter multi-view consistency, which prevents the addition of redundant Gaussians.  Tab. 7 further validates this by showing that with VCD, the number of Gaussians is reduced by 80%. This ablation study demonstrates the effectiveness of VCD for accelerating training.
Multi-View Consistent Pruning. Next, we evaluate the effectiveness of VCP. As shown in  Tab. 7, adding VCP shortens the training time by 25% and reduces the number of Gaussians by 26%, without sacrificing rendering quality. This is because our method effectively removes redundant Gaussians while preserving those critical for scene reconstruction. The strict multi-view consistency evaluation for each deleted Gaussian ensures this effectiveness, demonstrating that VCP is highly effective.
Compact Box. Finally, we evaluate the effectiveness of compact box. As shown in  Tab. 7, adding CB shortens the training time by 14% while achieving comparable rendering quality. This demonstrates that CB can accelerate training without degrading reconstruction quality.
Table 6: Quantitative results of accelerating various backbones. Our method achieves a 3-8× training speed-up.
Method Mip-NeRF 360 [barron2022mip]
[…]
Mip-Splatting [yu2024mip] 26.20 27.89 0.837 0.176 3.98M 224
+Ours 2.98 27.95 0.828 0.208 0.83M 606
Scaffold-GS [lu2024scaffold] 18.37 27.70 0.812 0.226 0.57M 194
+Ours 5.06 27.68 0.809 0.220 0.30M 281
@@
3DGS-accel 7.10 27.46 0.810 0.226 2.64M 182
+VCD. 3.53 27.69 0.798 0.259 0.53M 222
+VCP. 5.32 27.70 0.812 0.228 1.96M 285
+CB. 6.13 27.44 0.810 0.223 2.78M 303
Full 1.93 27.56 0.797 0.261 0.40M 579